OpenStudy (anonymous):

(Calculus 1) Help! Prove that the following function is NOT bounded

OpenStudy (anonymous):

Function attached below

OpenStudy (anonymous):

oops correction: it's 2x sin(1/x^2)

OpenStudy (shamil98):

first off, what does not bounded mean?

OpenStudy (anonymous):

that f(x) can go as high and as low as possible

OpenStudy (shamil98):

so no restrictions right?

OpenStudy (anonymous):

a bounded function is |f(x)|<M ,M is the bound

OpenStudy (shamil98):

Okay, so basically you have to prove that it has no restrictions then.

OpenStudy (anonymous):

What do you mean by no restrictions?

OpenStudy (shamil98):

bounded = means there are restrictions, correct?

OpenStudy (anonymous):

I think you mean the domain of f(x)?

OpenStudy (shamil98):

no idea, i have no sleep on me xD

OpenStudy (shamil98):

@TuringTest Care to help with this? I'm a bit exhausted... >.<

OpenStudy (anonymous):

Heh, it's ok. I appreciate that you try to help :)

OpenStudy (turingtest):

just take the limit as x goes to infinity and -infinity

OpenStudy (turingtest):

\[f(x)=2x\sin(\frac1{x^2})-{\cos(\frac1{x^2})\over x^2}-\frac12\]

OpenStudy (anonymous):

Taking said limits would lead to an assumption that f(x) is bounded as the limits are both equal to 0

OpenStudy (anonymous):

The area where f(x) takes values close to infinity and negative infinity is near 0

OpenStudy (turingtest):

how is\[\lim_{x\to\infty} f(x)=0\]?

OpenStudy (turingtest):

that's not what the question is asking, they want to know if the funciton is bounded as x ranges from -infty to +inty

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

\[\lim_{x\to\infty}f(x)=\lim_{x\to\infty}2x\sin(\frac1{x^2})-\lim_{x\to\infty}{\cos(\frac1{x^2})\over x^2}-\lim_{x\to\infty}\frac12\]what is the value of each limit?

OpenStudy (anonymous):

its clear from the graph that the limits are going to 0

OpenStudy (turingtest):

yeah, except they don't

OpenStudy (anonymous):

okay they go to -1/2 and 1/2, but how does that prove anything?

OpenStudy (turingtest):

they don't do that either, what is\[\lim_{x\to\infty}2x\sin(\frac1{x^2})\]?

OpenStudy (anonymous):

0?

OpenStudy (turingtest):

right, and as \(x\to-\infty\)

OpenStudy (turingtest):

?

OpenStudy (anonymous):

same i guess

OpenStudy (turingtest):

yep, and the next term?

OpenStudy (anonymous):

to 0 as well

OpenStudy (turingtest):

right so we are left with... ?

OpenStudy (anonymous):

-1/2

OpenStudy (turingtest):

for both +/- infty, yes therefore the function is bounded both above and below by that number

OpenStudy (anonymous):

dude, i think you are mistaken. we can only predict the range of a function using limits only if its monotonous.

OpenStudy (turingtest):

oh I see, I did end behavior, so you want the range?

OpenStudy (anonymous):

yeah if we manage to show that the range is a not bounded then its proven

OpenStudy (turingtest):

then take the limit as x goes to 0, since that is where we are likely to find the maximum values of the function (when we have some 0's in the denom)

OpenStudy (anonymous):

hmmm, the limit as x approaches 0 doesnt exist as it fluctuates between -infinity and infinity

OpenStudy (anonymous):

does that mean anything?

OpenStudy (anonymous):

i mean can this be used as proof?

OpenStudy (turingtest):

yeah, it means the range is unbounded

OpenStudy (turingtest):

yes, showing that the left hand limit and right hand limit of the cosine term approach +/- infty means that the function has no upper nor lower bound.

OpenStudy (anonymous):

GREAT. THATS GOOD HELP :)

OpenStudy (turingtest):

sorry I sent you off-track there for a while

OpenStudy (anonymous):

no problem we got to the point and that helped me so much!